one watering system needs about 3 times as long to complete a job as another watering system.when both systems operate at the same time,the job can be completed in 9 mins.how long does it take each systemto do the job alone?
Answer by checkley71(8403)

XY/(X+Y) IS THE FORMULA.
LET Y=3X
X*3X/(X+3X)=9 CROSS MULTIPLY
3X^2=9(X+3X)
3X^2=9*4X
3X^2-36X=0
3X(X-12)=0
X-12=0
X=12 MINUTES FOR THE FASTER WATRERING SYSTEM.
3*12=36 MINUTES FOR THE SLOWER WATERING SYSTEM.
PROOF
12*36/(12+36)=9
432/48=9
9=9